OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −4(3d − 1) + 4d

OpenStudy (anonymous):

−4(3d − 1) + 4d =(-4)(3d)-4(-1)+4d =-12d+4+4d =-8d+4